# Multi-objective Reinforcement Learning with Nonlinear Preferences: Provable Approximation for Maximizing Expected Scalarized Return

### v3

@@ 5 Computing Optimal Policies
By contrast, a brute force algorithm that conditions on all previous trajectories, even with these assumptions must scale no better than $O(|\mathcal{S}|^{T})$.
6 Experiments and Empirical Observations
@@
We test RAVI on two distinct interpretations of multi-objective reinforcement learning: A. the fairness interpretation where the agent tries to maximize rewards across all dimensions. B. the objective interpretation, where the agent tries to maximize one while minimizing the other. In both scenarios, we show that RAVI can discover more optimal policies than other baselines for nonlinear multi-objective optimization. This holds even for coarser settings of $\alpha$ in the algorithm than would be necessary for strong theoretical worst-case approximation guarantees.
6.1 Taxi
@@
We use the taxi multi-objective simulation environment considered in [13] for testing nonlinear ESR maximization. In this grid-world environment (Figure 3), the agent is a taxi driver whose goal is to deliver passengers. There are multiple queues of passengers, each having a given pickup and drop-off location. Each queue is a different objective or dimension of the reward. The taxi can only take one passenger at a time and wants to balance trips for passengers in the different queues.
We adopt the Nash social welfare function $W_{\text{Nash}}(\mathbf{r})=\left(\prod_{i}r_{i}\right)^{1/d}$ and egalitarian welfare $W_{\text{Egalitarian}}(\mathbf{r})=\min\{r_{i}\}_{i}$ for fairness considerations. To ensure numerical stability, Nash Welfare is optimized with its smoothed log-transform $W_{\text{SPF}}(\mathbf{r},\lambda)=\sum_{i=1}^{d}\ln(r_{i}+\lambda)$ for $\lambda=10^{-7}$. During evaluations, we still use $W_{\text{Nash}}$.
We consider two intuitive baseline approaches for optimization in a resource allocation task like this one. A. Linearly Scalarized Policy [23]: A relatively straightforward technique for multi-objective RL optimization is to apply the linear combination on the Q-values for each objective. Given weights $\mathbf{w}\in\mathbb{R}^{d}$, $\sum_{i}^{d}w_{i}=1$, the scalarized objective is $SQ(s,a)=\mathbf{w}^{\top}\mathbf{Q}(s,a)$, where $Q(s,a)_{i}$ is the Q-value for the $i^{\text{th}}$ objective. B. Mixture Policy [22]: this baseline works by combining multiple Pareto optimal base policies into a single policy. The mixing is achieved by adopting a base policy for $I$ steps and then switching to the next.
All the hyperparameters, including the interval for mixture policy and the weights for linear scalarization, are tuned via grid search.
We run RAVI along with the baselines on 20 random initializations of a $10\times 10$ taxi environment with three passenger spawn locations and destinations ($d=3$), $T=100$ without reward discount ($\gamma=1$), and compare their scalarized total rewards at convergence. As shown in 1, we observe that RAVI is able to compute better performing policies across both nonlinear objectives compared to both baselines.
6.2 Scavenger Hunt
@@
Inspired by the Resource Gathering environment [8], the scavenger hunt environment is a grid-world simulation where the agent must collect resources while avoiding enemies scattered across the grid. The state representation includes the agent’s position and the status of the resources (collected or not). The reward function is vector-valued, where the first component indicates the number of resources collected, and the second component indicates the damage taken by encountering enemies.
Figure 2: Illustration of the axi environment. The taxi is represented by the yellow circle and passenger spawn and delivery locations are represented by colored cells. Note that the scale of the actual experiments are much larger.
Figure 3: Illustration of the scavenger hunt environment. The agent is represented by the blue cell, the green cells are resources, and the red cells are enemies.
Nash Welfare Egalitarian Welfare RAVI 5.06$\pm$0.13 4.35$\pm$0.49 Mixture Policy 3.99$\pm$0.43 2.86$\pm$0.33 Linearly Scalarized Policy 0.02$\pm$0.26 0.03$\pm$0.14
Table 1: Comparison with baselines in terms of ESR for the taxi environment.
Cobb-Douglas RD Threshold RAVI 1.63$\pm$0.20 3.12$\pm$0.48 Welfare Q-learning 1.044$\pm$0.20 1.33$\pm$0.54 Linearly Scalarized Policy 1.02$\pm$0.32 1.68$\pm$0.50
Table 2: Comparison with baselines in terms of ESR for the scavenger hunt environment.
Given the differing nature of the task from the taxi environment, where the objective is to maximize one reward component while minimizing the other, we adopt two specific scalarization functions for the scavenger hunt environment: A. Resource-Damage (RD) Threshold Scalarization: $RD_{\text{threshold}}(R,D)=R-\max(0,(D-\text{threshold})^{3})$, for the number of resources collected $R$ and the damage taken $D$. The threshold parameter, i.e. the budget, determines the point at which the penalty on the damage taken starts to apply. The RD threshold scalarization encodes the concept of a budget into the optimization process. It is a constrained optimization approach with a soft penalty on the damage taken. B. Cobb-Douglas Scalarization: $CD(R,D)=R^{\alpha}\left(\frac{1}{D+1}\right)^{\beta}$. The Cobb-Douglas scalarization function is inspired by economic theory. It balances the trade-off between the number of resources collected and the damage taken. The parameters $\alpha$ and $\beta$ represent the weights assigned to each component, with $\alpha+\beta=1$. Note that when $\alpha$ and $\beta$ are set to 0.5, this scalarization function reduces $W_{\text{Nash}}(R,1/(D+1))$.
Due to the differing nature of the scavenger hunt task, a mixture policy does poorly — sometimes it ignores the resources and just tries not to hit an enemy, other times it ignores the enemies and just tries to get resources. We therefore compare the performance of RAVI against a linearly scalarized policy and Welfare Q-learning [13]: by using nonlinear updates and a non-stationary policy, this algorithm adapts the Q-learning algorithm to optimize ESR for a potentially nonlinear welfare function.
Similarly, we run both RAVI and Welfare Q-learning on 20 random initializations of a $15\times 15$ scavenger hunt environment with six resources and 40% of the cells randomly populated with enemies, $T=20$ without reward discount ($\gamma=1$), and compare their scalarized total rewards at convergence. As shown in table 2, we again observe that RAVI is able to compute better performing policies across both nonlinear objectives compared to Welfare Q-learning.
Discretization of Accumulated Rewards.
@@
We include an ablation study on the setting of the discretization interval length $\alpha$ in the appendix, in wihch we test values of $\alpha$ are substantially greater (thus making the algorithm more runtime efficient) than what would be necessary for worst-case theoretical guarantees from the previous analysis. Note that the empirical performance of RAVI is substantially better than is guaranteed by the previous theoretical analysis, for which substantially smaller $\alpha$ values would be necessary for worst-case guarantees.
7 Removing Knowledge of the Model
@@
We have shown that RAVI can efficiently find an approximately optimal policy given access to a model of the environment. In this section we observe that the model can be jointly learned by extending the classical E3 algorithm [16] to the nonlinear multi-objective case by lifting the exploration algorithm to the multi-objective setting and using RAVI for the exploitation subroutine. We call the resulting combined algorithm Reward-Aware-Explore or Exploit (or RAEE for short). We briefly explain the high level ideas and state the main result here and defer further discussion to the appendix due to space constraints.
@@
The algorithm consists of two stages, exploration and exploitation. Repeat the following stages, described here at a high level and in the appendix in more detail.
@@
  • Explore. At a given state, choose the least experienced action and record the reward and transition. Continue in this fashion until reaching a known state, where we say a state is known if it has been visited sufficiently many times for us to have precise local statistics about the reward and transition functions at that state.
@@
  • Exploit. Run RAVI from the current known state $s$ on the induced model comprising the known states and with a single absorbing state representing all unknown states. If the welfare obtained by this policy is within the desired error bound of $V^{*}(s,\mathbf{0},T)$ then we are done. Otherwise, compute a policy to reach an unknown state as rapidly as possible and resume exploring.
@@
Let $V^{*}(s,0,T)$ denote the value function for the policy with the optimal expected welfare in the MOMDP $M$ starting at state $s$, with $\mathbf{0}\in\mathbb{R}^{d}$ accumulated reward and $T$ timesteps remaining. Then for a uniformly continuous welfare function $W$, there exists an algorithm $A$, taking inputs $\epsilon$, $\beta$, $\mathcal{S}$, $\mathcal{A}$, and $V^{*}(s,\mathbf{0},T)$, such that the total number of actions and computation time taken by $A$ is polynomial in $1/\epsilon$, $1/\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, the horizon time $T=1/(1-\gamma)$ and exponential in the number of objectives $d$, and with probability at least $1-\beta$, $A$ will halt in a state $s$, and output a policy $\hat{\pi}$, such that $V^{\hat{\pi}}_{M}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$.
We provide additional details comparing our analysis with that of [16] in the appendix.
8 Conclusion and Future Work
Nonlinear prefernces in reinforcement learning are important, as they can encode fairness as the nonlinear balancing of priorities across multiple objectives or risk attitudes with respect to even a single objective. Stationary policies are not necessarily optimal for such objectives. We derived an extended form of Bellman optimality to characterize the structure of optimal policies conditional on accumulated reward. We introduced the RAVI and RAEE algorithms for efficiently computing an approximately optimal policy. Our primary contribution is a theoretical analysis, but we have also supplemented with an experimental demonstration.

### v4

@@ 5. Computing Optimal Policies
This dependence is significantly better than a naive brute-force approach, which scales at least as $O\bigl{(}|\mathcal{S}|^{T}\bigr{)}$.
Intuitively, the key savings arise from discretizing the reward space (with granularity $\delta_{\epsilon}$) rather than enumerating all possible trajectories. This discretization is guided by the smoothness assumption on $W_{\text{SPF}}$.
6. Experiments
Scavenger
Taxi, $d=5$
Figure 2. Visualization of the Taxi and Scavenger environments.
@@ 6. Experiments
We test RAVI on two distinct interpretations of multi-objective reinforcement learning: (1) the fairness interpretation, where the agent tries to maximize rewards across all dimensions. (2) the objective interpretation, where the agent tries to maximize one while minimizing the other. In both scenarios, we show that RAVI can discover more optimal policies than other baselines for nonlinear multi-objective optimization. This holds even for coarser settings of $\alpha$ in the algorithm than would be necessary for strong theoretical worst-case approximation guarantees.
6.1. Simulation Environments
@@
Taxi: We use the taxi multi-objective simulation environment considered in Fan et al. (2023) for testing nonlinear ESR maximization. In this grid-world environment, the agent is a taxi driver whose goal is to deliver passengers. There are multiple queues of passengers, each having a given pickup and drop-off location. Each queue has a different objective or dimension of the reward. The taxi can only take one passenger at a time and wants to balance trips for passengers in the different queues. This environment models the fairness interpretation.
@@
Scavenger: Inspired by the Resource Gathering environment Barrett and Narayanan (2008), the scavenger hunt environment is a grid-world simulation where the agent must collect resources while avoiding enemies scattered across the grid. The state representation includes the agent’s position and the status of the resources (collected or not). The reward function is vector-valued, where the first component indicates the number of resources collected, and the second component indicates the damage taken by encountering enemies. This environment is the objective interpretation.
6.2. Baseline Algorithms
Linearly Scalarized Policy (LinScal) Van Moffaert et al. (2013): A relatively straightforward technique for multi-objective RL optimization is to apply the linear combination on the Q-values for each objective. Given weights $\mathbf{w}\in\mathbb{R}^{d}$, $\sum_{i}^{d}w_{i}=1$, the scalarized objective is $SQ(s,a)=\mathbf{w}^{\top}\mathbf{Q}(s,a)$, where $Q(s,a)_{i}$ is the Q-value for the $i^{\text{th}}$ objective. $\epsilon$-greedy policy is used on $SQ(s,a)$ during action selection, and regular Q-learning updates are applied on each reward dimension.
Mixture Policy (Mixture) Vamplew et al. (2009): this baseline works by combining multiple Pareto optimal base policies into a single policy. Q-learning is used to optimize for each reward dimension separately (which is approximately Pareto optimal), and the close-to-optimal policy for each dimension is used for $I$ steps before switching to the next.
Welfare Q-learning (WelfareQ) Fan et al. (2023): this baseline extends Q-learning in tabular setting to approximately solve the nonlinear objective function by considering past accumulated rewards to perform non-stationary action selection.
Model-Based Mixture Policy (Mixture-M): Instead of using Q-learning to find an approximately Pareto optimal policy for each objective, value iteration Sutton et al. (1998) is used to calculate the optimal policy, and each dimension uses the corresponding optimal policy for $I$ steps before switching to the next.
6.3. Nonlinear Functions
Taxi: We use the following three functions on Taxi for fairness considerations: (1) Nash social welfare: $W_{\text{Nash}}(\mathbf{r})=\left(\prod_{i}r_{i}\right)^{1/d}$. (2) Egalitarian welfare: $W_{\text{Egalitarian}}(\mathbf{r})=\min\{r_{i}\}_{i}$. (3) $p$-welfare§§§$p$-welfare is equivalent to generalized mean. If $p\rightarrow 0$, $p$-welfare converges to Nash welfare; if $p\rightarrow-\infty$, $p$-welfare converges to egalitarian welfare; and $p$-welfare is the utilitarian welfare when $p=1$.: $W_{p}(\mathbf{r})=(\frac{1}{d}\sum_{i}r_{i}^{p})^{1/p}$.
Scavenger: We use these two functions to reflect the conflicting nature of the objectives:
  1. (1)
    Resource-Damage Threshold Scalarization: $RD_{\text{threshold}}(R,D)=R-\max(0,(D-\text{threshold})^{3})$, where $R$ is number of resources collected and $D$ is damage taken from enemies. The threshold parameter represents a budget after which the penalty from the damage starts to apply.
  2. (2)
    Cobb-Douglas Scalarization: $CD_{\rho}(R,D)=R^{\rho}\left(1/(D+1)\right)^{(1-\rho)}$. This function is inspired by economic theory and balances trade-offs between $R$ and $D$.
6.4. Experiment Settings and Hyperparameters
We run all the algorithms using 10 random initializations with a fixed seed each. We set $\mathbf{w}=(1/d)\times\mathbf{1}$ (uniform weight) for LinScal, $I=T/d$ for Mixture and Mixture-M, $\alpha=1$ for RAVI, and learning rate of 0.1 for WelfareQ. Three of our baselines (Mixture, LinScal, and WelfareQ) are online algorithms. Thus, to ensure a fair comparison, we tuned their hyperparameters using grid-search and evaluated their performances after they reached convergence. For model-based approaches (RAVI and Mixture-M), we run the algorithms and evaluate them after completion. We set the convergence threshold to $\Delta=10^{-7}$ for Mixture-M. Some environment-specific settings are discussed below.
Taxi: To ensure numerical stability of $W_{\text{Nash}}$, we optimize its smoothed log-transform $W_{\text{SPF}}(\mathbf{r},\lambda)=\sum_{i=1}^{d}\ln(r_{i}+\lambda)$, but during evaluations we still use $W_{\text{Nash}}$. We set $\lambda=10^{-8}$, $T=100$. $\gamma=1$, size of the grid world to be $15\times 15$ with $d\in\{2,3,4,5\}$ reward dimensions.
Scavenger: We set threshold for $RD_{\text{threshold}}$ to 2 and $\rho=0.4$ for $CD_{\rho}$. The size of the grid world is $15\times 15$ with six resources scattered randomly and 1/3 of the cells randomly populated with enemies. We use $T=20$ and $\gamma=1$.
Table 1. Comparison with baselines in terms of ESR.
Environment Dimension Function RAVI Mixture LinScal WelfareQ Mixture-M
Taxi $d=2$ $W_{\text{Nash}}$ 7.555$\pm$0.502 5.136$\pm$0.547 0.000$\pm$0.000 5.343$\pm$0.964 6.406$\pm$0.628
$W_{\text{Egalitarian}}$ 3.000$\pm$0.000 3.483$\pm$0.894 0.000$\pm$0.000 0.387$\pm$0.475 4.065$\pm$0.773
$W_{p=-10}$ 5.279$\pm$0.000 3.623$\pm$0.522 0.000$\pm$0.000 2.441$\pm$1.518 4.356$\pm$0.829
$W_{p=0.001}$ 7.404$\pm$0.448 5.363$\pm$0.844 0.000$\pm$0.000 4.977$\pm$1.514 6.406$\pm$0.628
$W_{p=0.9}$ 9.628$\pm$0.349 5.947$\pm$0.488 7.833$\pm$0.908 7.999$\pm$0.822 7.052$\pm$0.412
$d=3$ $W_{\text{Nash}}$ 4.996$\pm$0.297 3.250$\pm$0.584 0.000$\pm$0.000 2.798$\pm$1.462 3.461$\pm$0.459
$W_{\text{Egalitarian}}$ 2.000$\pm$0.000 2.030$\pm$0.981 0.000$\pm$0.000 0.094$\pm$0.281 1.660$\pm$0.663
$W_{p=-10}$ 3.115$\pm$0.000 2.129$\pm$0.878 0.000$\pm$0.000 2.558$\pm$0.812 1.849$\pm$0.733
$W_{p=0.001}$ 3.307$\pm$0.000 3.118$\pm$0.536 0.000$\pm$0.000 3.306$\pm$1.181 3.462$\pm$0.458
$W_{p=0.9}$ 6.250$\pm$0.322 3.883$\pm$0.329 5.191$\pm$0.464 4.834$\pm$0.797 4.081$\pm$0.266
$d=4$ $W_{\text{Nash}}$ 2.191$\pm$0.147 0.579$\pm$0.713 0.000$\pm$0.000 1.601$\pm$0.189 1.122$\pm$0.781
$W_{\text{Egalitarian}}$ 1.700$\pm$0.483 0.545$\pm$0.445 0.000$\pm$0.000 0.000$\pm$0.000 0.634$\pm$0.437
$W_{p=-10}$ 1.029$\pm$0.000 0.490$\pm$0.490 0.000$\pm$0.000 0.689$\pm$0.451 0.688$\pm$0.475
$W_{p=0.001}$ 2.145$\pm$0.129 0.961$\pm$0.628 0.000$\pm$0.000 1.440$\pm$0.511 1.126$\pm$0.774
$W_{p=0.9}$ 3.369$\pm$0.173 1.230$\pm$0.264 2.546$\pm$0.174 2.521$\pm$0.211 1.689$\pm$0.320
$d=5$ $W_{\text{Nash}}$ 2.308$\pm$0.185 0.000$\pm$0.000 0.000$\pm$0.000 1.181$\pm$0.613 0.000$\pm$0.000
$W_{\text{Egalitarian}}$ 1.700$\pm$0.483 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000
$W_{p=-10}$ 1.023$\pm$0.000 0.000$\pm$0.000 0.000$\pm$0.000 0.407$\pm$0.501 0.000$\pm$0.000
$W_{p=0.001}$ 2.000$\pm$0.102 0.018$\pm$0.020 0.000$\pm$0.000 1.309$\pm$0.490 0.025$\pm$0.025
$W_{p=0.9}$ 3.289$\pm$0.147 1.220$\pm$0.141 2.844$\pm$0.250 2.879$\pm$0.263 1.761$\pm$0.148
Scavenger $d=2$ $CD_{\rho=0.4}$ 1.336$\pm$0.240 0.655$\pm$0.558 0.874$\pm$0.605 0.713$\pm$0.645 0.729$\pm$0.296
$RD_{\text{threshold}=2}$ 3.400$\pm$0.843 0.900$\pm$0.943 1.200$\pm$0.872 1.444$\pm$1.423 0.845$\pm$2.938
Taxi, $W_{\text{Nash}}$, $d=5$
Taxi, $W_{p=0.9}$, $d=4$
Scavenger, $CD_{\rho=0.4}$
Scavenger, $RD_{\text{threshold}}$
Taxi, $W_{\text{Nash}}$, $d=2$
Taxi, $W_{p=0.9}$, $d=3$
Taxi, $W_{p=0.001}$, $d=4$
Taxi, $W_{p=-10}$, $d=5$
Figure 3. Comparisons with baselines, with learning curves included.
6.5. Results
As shown in Table 1, we found that RAVI is able to generally outperform all baselines across all settings in both Taxi and Scavenger environments in terms of optimizing our nonlinear functions of interest.
On the Taxi environment, we observe that LinScal is unable to achieve any performance except for $W_{p=0.9}$. This is an expected behavior due to the use of a linearly scalarized policy and the fact that $p$-welfare converges to utilitarian welfare when $p\rightarrow 1$. Among the baseline algorithms, we observe that WelfareQ performs the best on $d=5$, whereas Mixture and Mixture-M fail due to the use of a fixed interval length for each optimal policy. Furthermore, the advantage of RAVI becomes more obvious as $d$ increases.
On the Scavenger environment, due to $d=2$ and having a dense reward signal, all algorithms are able to achieve reasonable values for the welfare functions. We also observe that RAVI significantly outperforms all baselines.
Given that Mixture, LinScal, and WelfareQ are online algorithms, for the sake of completeness, we provide the visualizations of the learning curves of the baseline algorithms compared with RAVI with different $\alpha$ values in Figure 3, where the model-base approaches are shown as horizontal dotted lines. In general, we found that the online algorithms tend to converge early, and we observe a degradation in performance as $\alpha$ increases. The full set of results can be found at Section C.3.
6.6. Ablation Study on Discretization Factor
To further evaluate the effect of $\alpha$ on RAVI, we also run two sets of experiments:
  1. (1)
@@ 6.6. Ablation Study on Discretization Factor
    we test RAVI performance on settings with $\gamma<1$ and adopt $\alpha$ values that are substantially greater than those necessary for worst-case theoretical guarantees from the previous analysis. This set of results can be found at Section C.1. Note that the empirical performance of RAVI is substantially better than is guaranteed by the previous theoretical analysis.
  2. (2)
    With $\gamma=1$, we evaluate RAVI using larger alpha values and investigate how much performance degradation occurs. This set of experiments can be found in Section C.2.
7. Removing Knowledge of the Model
@@
We have shown that RAVI can efficiently find an approximately optimal policy given access to a model of the environment. In this section we observe that the model can be jointly learned by extending the classical E3 algorithm Kearns and Singh (2002) to the nonlinear multi-objective case by lifting the exploration algorithm to the multi-objective setting and using RAVI for the exploitation subroutine. We call the resulting combined algorithm Reward-Aware-Explore or Exploit (or RAEE for short). We briefly explain the high level ideas and state the main result here and defer further discussion to the appendix due to space constraints.
@@
The algorithm consists of two stages, exploration and exploitation. The algorithm alternates between two stages: exploration and exploitation. Each stage is outlined here at a high level, with more detailed steps provided in the appendix.
@@
  • Explore. At a given state, choose the least experienced action and record the reward and transition. Continue in this fashion until reaching a known state, where we say a state is known if it has been visited sufficiently many times for us to have precise local statistics about the reward and transition functions in that state.
@@
  • Exploit. Run RAVI from the current known state $s$ in the induced model comprising the known states and with a single absorbing state representing all unknown states. If the welfare obtained by this policy is within the desired error bound of $V^{*}(s,\mathbf{0},T)$, then we are done. Otherwise, compute a policy to reach an unknown state as quickly as possible and resume exploring.
@@
Let $V^{*}(s,0,T)$ denote the value function of the policy with the optimal expected welfare in the MOMDP $M$ starting at state $s$, with $\mathbf{0}\in\mathbb{R}^{d}$ accumulated reward and $T$ timesteps remaining. Then for a uniformly continuous welfare function $W$, there exists an algorithm $A$, taking inputs $\epsilon$, $\beta$, $\mathcal{S}$, $\mathcal{A}$, and $V^{*}(s,\mathbf{0},T)$, such that the total number of actions and computation time taken by $A$ is polynomial in $1/\epsilon$, $1/\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, the horizon time $T=1/(1-\gamma)$ and exponential in the number of objectives $d$, and with probability at least $1-\beta$, $A$ will halt in a state $s$, and output a policy $\hat{\pi}$, such that $V^{\hat{\pi}}_{M}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$.
We provide additional details comparing our analysis with that of Kearns and Singh (2002) in the appendix.
As we do not regard the learning of the Multi-Objective Markov Decision Process (MOMDP) as our primary contribution, we choose to focus the empirical evaluation on the nonlinear optimization subroutine, which is the most crucial modification from the learning problem with a single objective. The environments we used for testing the optimality of RAVI are very interesting for optimizing a nonlinear function of the rewards, but are deterministic in terms of transitions and rewards, making the learning of these environments less interesting empirically.
8. Conclusion and Future Work
Nonlinear preferences in reinforcement learning are important, as they can encode fairness as the nonlinear balancing of priorities across multiple objectives or risk attitudes with respect to even a single objective. Stationary policies are not necessarily optimal for such objectives. We derived an extended form of Bellman optimality to characterize the structure of optimal policies conditional on accumulated reward. We introduced the RAVI and RAEE algorithms to efficiently compute an approximately optimal policy.
Our work is certainly not the first to study MORL including with nonlinear preferences. However, to the best of our knowledge, our work is among the first to provide worst-case approximation guarantees for optimizing ESR in MORL with nonlinear scalarization functions.
While our experiments demonstrate the utility of RAVI in specific settings, there are many possible areas of further empirical evaluation including stochastic environments and model learning alongside the use of RAVI as an exploitation subprocedure as described in theory in Section 7 with RAEE. Further details on the limitations of our experiments are discussed in Appendix E.